Two angles are complementary. The second angle is 30 degrees more than the first. Write an equation and solve the measures of both angles.
Found 2 solutions by jorel555, Alan3354:
Answer by jorel555(1290)

If the angles are complementary, they add up to 90 degrees. So:
X+(X+30)=90
2X+30=90
2X=60
X=30
X+30=60
The angles are 30 and 60 degrees, respectively!!!!!!!!!!!!

Answer by Alan3354(69443)

Two angles are complementary. The second angle is 30 degrees more than the first. Write an equation and solve the measures of both angles.
--------------
Angles are A and B
---
A + B = 90 degs
A = .
B + 30